# Beyond the Answers: Reviewing the Rationality of Multiple Choice Question Answering for the Evaluation of Large Language Models

### v1

@@ Abstract
In the field of natural language processing (NLP), Large Language Models (LLMs) have precipitated a paradigm shift, markedly enhancing performance in natural language generation tasks. Despite these advancements, the comprehensive evaluation of LLMs remains an inevitable challenge for the community. Recently, the utilization of Multiple Choice Question Answering (MCQA) as a benchmark for LLMs has gained considerable traction. This study investigates the rationality of MCQA as an evaluation method for LLMs. If LLMs genuinely understand the semantics of questions, their performance should exhibit consistency across the varied configurations derived from the same questions. Contrary to this expectation, our empirical findings suggest a notable disparity in the consistency of LLM responses, which we define as REsponse VAriability Syndrome (REVAS) of the LLMs, indicating that current MCQA-based benchmarks may not adequately capture the true capabilities of LLMs, which underscores the need for more robust evaluation mechanisms in assessing the performance of LLMs.
Machine Learning, ICML
Research Center for Social Computing and Information Retrieval, Harbin Institute of Technology, China
{hcwang,sdzhao@ir.hit.edu.cn}
@@ 1 Introduction
The emergence of Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023a), Baichuan (Yang et al., 2023), and ChatGPT (OpenAI, 2022), represents a paradigm shift in the field of Natural Language Processing (NLP). These models have exhibited exceptional proficiency in mimicking human-like textual outputs, establishing their significance across a diverse array of applications. However, the challenge of effectively evaluating LLMs persists (Chang et al., 2023). This difficulty arises from the intricate nature of natural language. Conventional evaluation metrics for generative tasks often fall short in accurately assessing the performance of LLMs, since most LLMs can generate text contextually rich and coherent (Thoppilan et al., 2022), complicating the assessment of the outputs through merely quantitative evaluation based on text matching such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004).
@@
Multiple-Choice Question Answering (MCQA) is a fundamental format for various tasks in NLP, such as common sense reasoning (Talmor et al., 2019; Sap et al., 2019; Zellers et al., 2018), reading comprehension (Lai et al., 2017; Huang et al., 2019) and cloze-style tasks (Zellers et al., 2019; Mostafazadeh et al., 2016). Each MCQA instance comprises a question paired with several answer options, requiring models to identify the correct response as depicted in Figure 1. As a non-subjective metric, MCQA serves as a prominent automatic evaluation method with accuracy as evaluation metric for numerous LLMs to test for the commonsense knowledge or knowledge for specific domain (Gao et al., 2021; Touvron et al., 2023a; OpenAI et al., 2023).
Figure 1: An example of Multiple-Choice Question Answering and two ranking strategies.
@@
Despite the advanced performance of LLMs on the accuracy of MCQA-format benchmarks like MMLU (Hendrycks et al., 2021), a key challenge that persists in evaluating LLMs is maintaining consistency in responses when confronted with variations in answer options, and discrepancies arise even when the same questions are reformulated into the others (Robinson & Wingate, 2022; Wang et al., 2023; Zheng et al., 2023). The prior studies have revealed that LLMs exhibit sensitivity to the order of answer choices, leading to inconsistent responses, which underscores an issue that the accuracy of MCQA-format tasks may not reflect the authentic capability of LLMs. However, the fluctuation in the LLM responses due to the order of answer options may not be the only hurdle in the evaluation of LLMs.
In this study, we undertake a comprehensive examination through experiments involving various configurations based on the MCQA-format datasets: (1) original MCQA datasets, (2) modified MCQA datasets with altered answer options, and (3) conversion of MCQA-format instances into true/false format. In conducting this comprehensive evaluation, we deepen the insight into the limitations of LLMs in terms of response consistency. We define the phenomenon as REsponse VAriability Syndrome (REVAS), where the LLMs show inconsistencies in the questions derived from one question. Empirical results have proved that the high performance of the MCQA-format benchmark does not denote that the LLMs fully understand the semantics of the questions or learn the knowledge about the questions. Additionally, we explore the underlying reasons for observed behavioral patterns in LLMs through substantial experiments, that LLMs probably choose the most correct from the answer options and meanwhile regard some other options as correct as well. Therefore, we propose complementary training with negative samples to refine the LLMs, which could pave the way for developing more robust and adaptable NLP models, ultimately bridging the gap between machine and human-like understanding and reasoning in NLP tasks.
@@
  • We challenge the rationality of evaluating LLMs using MCQA-format benchmarks by involving various instances derived from the MCQA tasks.
  • Empirical findings suggest that LLMs may acquire a limited understanding of the MCQA-format questions, exhibiting limited generalization even with minor variations.
  • We hypothesize underlying causes for the observed patterns in LLMs and suggest potential strategies for enhancing the robustness and adaptability of LLMs.
2 Related Work
Transformer-based language models (Vaswani et al., 2017) have led the research of nature language processing (NLP) into a new era with the structure of self-attention and massive pre-training on large amounts of self-supervised corpora. Prior work has elucidated the scaling law, positing that the efficacy of language models is typically proportional to both the model size and the magnitude of pre-training data (Henighan et al., 2020; Kaplan et al., 2020). This principle has catalyzed a paradigm shift towards Large Language Models (LLMs), with the number of parameters escalating from millions (Devlin et al., 2019; Radford et al., 2019) to billions (Radford et al., 2018; Brown et al., 2020).
Recent advancements, including supervised fine-tuning and alignment with human values (Ouyang et al., 2022; Chung et al., 2022; Bai et al., 2022), have further augmented the capabilities of LLMs, enabling them to adhere more closely to human instructions and ethical considerations. Nonetheless, challenges persist since LLMs may show inconsistency in the model responses, especially under the scenarios of MCQA. Robinson & Wingate 2022 termed the ability to associate the answer options and corresponding symbols as multiple choice symbol binding (MCSB) and proved that the MCSB ability varied significantly by models. Additionally, Wang et al. 2023 revealed vulnerabilities in the ranking of candidate responses, which could be manipulated by altering the presentation order. Zheng et al. 2023 investigated the token bias in LLMs, which could lead to selection biases during option prediction. Furthermore, Kadavath et al. 2022 explored the reliability of the LLM performance and calibration, focusing exclusively on a set of private models under the MCQA settings. Their methodology involved modifying the options and transforming the original questions into binary-choice questions answering with options of “True” and “False”. Building on these foundational studies, in this study we conduct a comprehensive range of experiments applied to a diverse series of LLMs, including both open-source and commercial models. We aim to assess their performance consistency across various settings while using the same set of instances, thereby providing a broader understanding of their capabilities and limitations in different application scenarios.
@@ 2 Related Work
3 Can the Accuracy of MCQA-format Tasks Reflect the True Model Capability?
3.1 Overview
The assessment of LLMs has increasingly employed MCQA as a standardized metric. This method, featuring a single question accompanied by multiple answer choices, serves to gauge the performance of an LLM performance in specific tasks or domains. Notable MCQA datasets utilized in evaluating LLMs encompass a diverse range including multitask language comprehension (e.g., MMLU (Hendrycks et al., 2021)), benchmark assessments for LLMs (e.g., Big Bench (Srivastava et al., 2023)), commonsense reasoning (e.g., HellaSwag (Zellers et al., 2019), ARC (Clark et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021)), and human-centric evaluations (e.g., AGI-eval (Zhong et al., 2023)). Table 1 lists prominent LLMs such as Chinchilla (Hoffmann et al., 2022), Falcon (Almazrouei et al., 2023), FLAN (Chung et al., 2022), Galactica (Taylor et al., 2022), GPT4 (OpenAI et al., 2023), LLaMA2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and PaLM2 (Anil et al., 2023), along with their respective evaluations with MCQA-format tasks. Additionally, MCQA tasks extend to tests like AP exams for GPT4 and medical-specific MultiMedQA for Med-PaLM (Singhal et al., 2023). The adoption of MCQA tasks as evaluation of LLMs is based on the hypothesis that MCQA can effectively and accurately reveal the true capabilities of the LLMs. If so, LLMs must demonstrate an understanding of the contextual content and possess the requisite knowledge for correctly answering questions, thereby exhibiting consistencies across similar questions. To examine the above hypothesis, we conduct a series of experiments using configurations derived from several MCQA datasets, aimed at observing the behavior of the LLMs.
3.2 Experimental Design
In our study, we select three widely recognized LLMs: LLaMA2 (Touvron et al., 2023b), Baichuan2 (Yang et al., 2023), and ChatGPT 3.5 (OpenAI, 2022), for evaluation on the MMLU (Hendrycks et al., 2021) benchmark and the MedMCQA (Pal et al., 2022) dataset. MMLU represents a prevalent benchmark in MCQA for LLMs as presented in Table 1, while MedMCQA demands extensive medical knowledge and is challenging for most LLMs. The selected evaluation samples comprise 452 instances from MMLU and 967 instances from MedMCQA. These instances are specifically chosen for their adaptability to alternative formats, allowing for the creation of derived datasets. These derived datasets, encompassing a magnified sample size—37 times that of the initially selected instances—facilitate a more extensive and nuanced analysis of LLM capabilities, culminating in over 52,000 testing instances. Further details regarding the selected and constructed instances are available in the supplementary materials. Each experiment is conducted in both zero-shot and few-shot settings.
3.3 Adaptation of MCQA Answer Options
@@
Our primary method involves subtle modifications to the answer options within MCQA instances (as depicted in Figure 2). Ideally, an intelligent LLM should display resilience to such modifications, since the knowledge required by the question remains unchanged.
@@
Table 2: Distribution of $S_{c}$ (and corresponding accuracy) on MMLU with the original answer options under the 0-shot setting. %$S_{c}=n/4$ indicates the proportion of instances where LLMs consistently predicted $n$ out of 4 questions with re-ordered options. N/A: not applicable.
@@
%$S_{c}=4/4$ 43.02 (62.74) 19.03 (67.39) 48.23 (81.65) 34.54 (61.38) 52.43 (85.23) 40.43 (85.17)
%$S_{c}=3/4$ 34.73 (77.71) 37.83 (64.33) 38.27 (44.51) 40.74 (42.13) 35.17 (40.93) 39.81 (43.48)
%$S_{c}=2/4$ 20.35 (21.02) 29.37 (56.52) 11.28 (35.29) 17.89 (31.79) 11.50 (7.59) 17.89 (14.07)
%$S_{c}=1/4$ 7.08 (N/A) 8.58 (N/A) 2.21 (N/A) 6.83 (N/A) 0.88 (N/A) 1.86 (N/A)
Table 3: Performance deterioration in the LLMs post “None of the Above” substitution. Acc denotes the accuracy of LLMs with the initial order of the other options. Acc-H, termed as accuracy-hard, implies that an instance is deemed correctly predicted only if the LLMs demonstrate consistency across every permutation of the answer options. Red font delineates the performance decrement compared to that with the original options.

### v2

@@ Abstract
In the field of natural language processing (NLP), Large Language Models (LLMs) have precipitated a paradigm shift, markedly enhancing performance in natural language generation tasks. Despite these advancements, the comprehensive evaluation of LLMs remains an inevitable challenge for the community. Recently, the utilization of Multiple Choice Question Answering (MCQA) as a benchmark for LLMs has gained considerable traction. This study first investigates the limitations of MCQA as an evaluation method for LLMs and then analyzes the fundamental reason for the limitations of MCQA, that while LLMs may select the correct answers, it is possible that they also recognize other wrong options as correct. Finally, we propose a dataset augmenting method for Multiple-Choice Questions (MCQs), MCQA+, that can more accurately reflect the performance of the model, which underscores the need for more robust evaluation mechanisms in assessing the performance of LLMs.
@@ 1 Introduction
The emergence of Large Language Models (LLMs), such as GPT-3 [gpt3], LLaMA [touvron2023llama], and ChatGPT [chatgpt], represents a paradigm shift in the field of Natural Language Processing (NLP). These models have exhibited exceptional proficiency in mimicking human-like textual outputs, establishing their significance across a diverse array of applications. However, the challenge of effectively evaluating LLMs persists [chang2023survey]. This difficulty arises from the intricate nature of natural language. Conventional evaluation metrics for generative tasks often fall short in accurately assessing the performance of LLMs, since most LLMs can generate text contextually rich and coherent [thoppilan2022lamda], complicating the assessment of the outputs through merely quantitative evaluation based on text matching such as BLEU [papineni-etal-2002-bleu] and ROUGE [lin-2004-rouge].
@@
Multiple-Choice Question Answering (MCQA) is a fundamental format for various tasks in NLP, such as common sense reasoning [csqa, siqa, swag], reading comprehension [race, cqa] and cloze-style tasks [hs, sc]. Each MCQA instance comprises a question paired with several answer options, requiring models to identify the correct response as depicted in Figure 1. As a non-subjective metric, MCQA serves as a prominent automatic evaluation method with accuracy as an evaluation metric for numerous LLMs to test for the commonsense knowledge or knowledge for specific domain [leo_gao_2021_5371629, touvron2023llama, openai2023gpt4].
Figure 1: An example of MCQA and two ranking strategies.
@@
Despite the advanced performance of LLMs on the accuracy of MCQA-format benchmarks like MMLU [mmlu], previous studies have discussed a key challenge that persists in evaluating LLMs is maintaining invariability in responses when confronted with different orders of answer choices for a same question [robinson2022leveraging, wang2023large, zheng2023large], which underscores an issue that the accuracy of MCQA-format tasks may not reflect the authentic capability of LLMs. However, the above phenomenon may not be the only issue in the evaluation of LLMs with MCQA-format questions.
In this study, we conduct a detailed exploration through a series of experiments utilizing various configurations derived from original MCQA-format datasets. We aim to deepen the understanding of the inherent limitations of MCQA tasks as tools for evaluating LLMs with a focus on response invariability. Afterward, we explore the underlying reasons for the observed behavioral patterns in LLMs through substantial experiments, that LLMs probably choose the most correct from the answer options and meanwhile regard some other options as correct as well. Thus, evaluating LLM performance based on MCQA can yield ambiguous results. To address these challenges, we introduce an augmentation of the MCQA dataset, termed MCQA+, which includes variations of the original MCQs and is designed to more accurately reflect the true capabilities of LLMs. Empirical results demonstrate that the performance of LLMs on the MCQA+ dataset is significantly inferior to that on the original MCQA dataset, suggesting that our MCQA+ can facilitate the development of more robust and adaptable NLP models. This, in turn, may help narrow the gap between machine and human-like understanding and reasoning in NLP tasks.
@@
  • We challenge the rationality of using MCQA benchmarks for evaluating LLMs by substantial experiments.
  • We elucidate that within the MCQA context, LLMs may recognize multiple options as correct, but they ultimately select the option they assess as the most accurate.
  • We propose a dataset augmentation method that expands the original MCQA into MCQA+, which better reveals the models’ capacities.
@@
2 Limitations of MCQA Task as Benchmark for LLMs
2.1 MCQA Tasks as Benchmarks for LLMs
The assessment of LLMs has increasingly employed MCQA as a standardized metric. This method, featuring a single question accompanied by multiple answer choices, serves to gauge the performance of an LLM performance in specific tasks or domains. Notable MCQA datasets utilized in evaluating LLMs encompass a diverse range including multitask language comprehension (e.g., MMLU [mmlu]), benchmark assessments for LLMs (e.g., Big Bench [srivastava2023beyond]), commonsense reasoning (e.g., HellaSwag [hs], ARC [arc], PIQA [piqa], WinoGrande [w]), and human-centric evaluations (e.g., AGI-eval [zhong2023agieval]). Table 1 lists prominent LLMs such as Chinchilla [chinchilla], Falcon [almazrouei2023falcon], FLAN [chung2022scaling], Galactica [taylor2022galactica], GPT4 [openai2023gpt4], LLaMA2 [touvron2023llama2], Mistral [jiang2023mistral], and PaLM2 [anil2023palm2], along with their respective evaluations with MCQA-format tasks. Additionally, MCQA tasks extend to tests like AP exams for GPT4 and medical-specific MultiMedQA for Med-PaLM [singhal2023large]. The adoption of MCQA tasks as an evaluation of LLMs is based on the hypothesis that MCQA can effectively and accurately reveal the true capabilities of the LLMs. If so, LLMs must demonstrate an understanding of the contextual content and grasp the requisite knowledge for correctly answering questions, thereby exhibiting invariability across similar questions.
In this study, we select three widely recognized LLMs: LLaMA2 [touvron2023llama2], Baichuan2 [yang2023baichuan], and ChatGPT 3.5 [chatgpt], for evaluation on the MMLU [mmlu] benchmark and the MedMCQA [medmcqa] dataset. MMLU represents a prevalent benchmark in MCQA for LLMs as presented in Table 1, while MedMCQA demands extensive medical knowledge and is challenging for most LLMs. We adopt the greedy strategy during decoding to avoid the influence of token sampling. We conduct comprehensive experiments to demonstrate that the accuracy of LLMs on the MCQA tasks does not accurately reflect their capabilities concerning the specific knowledge required by the questions.
@@ 2.1 MCQA Tasks as Benchmarks for LLMs
2.2 Further Exploration with Re-ordered Answer Options
Prior research [robinson2022leveraging] has demonstrated that re-ordering answer options can lead to variability, where LLMs generate variant answers with the re-ordered answer options for the same question. In our study, we extend this exploration to include MMLU and MedMCQA datasets tested on LLaMA2, Baichuan2, and ChatGPT models. Our findings also corroborate the existence of such variability (details in Appendix A). Meanwhile, we prompt a deeper examination from two angles: (1) the impact of varying the number of answer options on performance variability; and (2) the correlation between this variability and the accuracy of model responses.
2.2.1 Variation in Answer Option Quantity
We explore the impact of varying the number of answer choices in MCQA tasks. Generally, MCQA-format tests feature four options per question [haladyna1993many]. We extend this framework by modifying the number of answer options to assess the adaptability of LLMs under these varied conditions.
When incrementing the number of answer options, we incorporate additional non-semantic tokens as distracting options to circumvent the unintentional introduction of plausible correct responses, thereby preserving the authenticity of the questions, as illustrated in Figure 2. This expands the number of answer options to 6, 8, and 10. In contrast, when reducing the number of options, we randomly eliminate distracting options, simplifying the questions by reducing the options to 3 or 2.
Figure 3: Performance of LLMs on MCQA tasks of MMLU and MedMCQA with variations in the number of options. “Accuracy” denotes the accuracy reported with the initial answer option order. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all corresponding numbers of re-ordered instances of an original question. Marked values in the sub-figures have 4 original options under the 0-shot setting.
By combining the adaptions of re-ordering and quantity variation of options, Figure 3 demonstrates the performance of LLMs on the MMLU and MedMCQA tasks. The metrics “Accuracy” and “Acc-H” are the same defined as in the preceding section. Intuitively, with fewer answer options, LLMs are expected to exhibit improved accuracy (achieving better “Accuracy” and maintaining “Acc-H”) as we have reduced the number of distracting options. Conversely, with increased answer options, LLMs are anticipated to maintain their accuracy since only non-distracting tokens (non-semantic tokens) have been added.
However, taking LLaMA2-13B as an illustrative example, it exhibits an accuracy of 72.73% with the original set of four options on MMLU under the zero-shot setting. However, with 2 to 10 answer options (other than 4 options), a marked volatility in accuracy is noted, ranging from 28.76% to 49.06%. Meanwhile, Acc-H exhibits a general declining trend as the number of options increases. This pattern of fluctuating accuracy and diminishing Acc-H values is consistent across all three LLMs on the MMLU. In experiments involving MedMCQA, a similar downward trend is observed for Accuracy, and Acc-H with the increment in answer options. Remarkably, for LLaMA2-13B, Acc-H plummets to 0.00% with six or more answer options, casting doubts on the reliability of the LLMs with a larger array of answer options. These phenomena underscore the unreliability of the MCQA task when the number of answer options is altered, as the following observation:
Observation 1: LLMs exhibit “overfitting” to MCQs with exactly four options. Consequently, the performance on MCQs with fewer or more than four options is not readily extrapolatable.
2.2.2 Correlation between the Variability and Accuracy
First, we conduct experiments with MCQA in its original answer option, recording baseline accuracy as “Acc”. Second, for simplicity, we expand the answer options in each MCQA-format question into four random distinct orders to assess the variability of LLMs to option sequencing as an example shown in Figure 2. We propose a novel evaluation metric, termed accuracy-hard (“Acc-H”), which quantifies the percentage of instances where LLMs invariantly and correctly predict across all four re-ordered orders. For example, Table 2 demonstrates that although LLMs like LLaMA2 can achieve an accuracy of 72.73% on the original MMLU, the Acc-H dropping to 26.99% indicates that LLaMA generates many different results if provided with answer options in different answer option orders. Afterward, Table 2 further analyzes the invariability (consistency) of LLMs when faced with four re-ordered answer options under the zero-shot setting. “$c=3/4$” indicates the subset of instances where three out of four predictions generated by LLMs are identical. It is noticeable that most of the LLMs can show invariability of no less than $3/4$. Notably, a higher $c$ value correlates with higher accuracy in the LLM responses. For instance, the accuracy of ChatGPT on MMLU is 85.23% when $c=4/4$, but it dwindles to 40.93% at $c=3/4$ and further plummets to 7.59% at $c=2/4$. This observed trend, consistent across various settings, suggests that the invariability of model responses across re-ordered instances is indicative of the model accuracy, which can be formalized as follows:
Observation 2: When an LLM consistently selects the same response to an MCQ, irrespective of the order in the sequence of the answer options, there is a high probability that the selected answer is correct.
Furthermore, the demonstration of invariability by LLMs in specific MCQs prompts us to ask the following question:
Question 1: If an LLM exhibits invariability in certain MCQs, can it be inferred to have genuinely mastered the knowledge involved in the questions?
@@ 2.2.2 Correlation between the Variability and Accuracy
Table 2: Model variability on MMLU and MedMCQA datasets. “Acc” denotes the accuracy on the original dataset. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all four re-ordered instances of an original question. Distribution of consistency (and corresponding accuracy) with the original answer options. “%$c=n/4$” indicates the percentage of instances where LLMs consistently predicted $n$ out of 4 questions with re-ordered options. N/A: not applicable. All experiments are under the 0-shot setting.
@@
Acc 72.73 43.23 70.21 45.81 65.00 56.88
Acc-H 26.99 12.82 39.38 21.20 44.69 34.44
$c=4/4$ 43.02 (62.74) 19.03 (67.39) 48.23 (81.65) 34.54 (61.38) 52.43 (85.23) 40.43 (85.17)
$c=3/4$ 32.69 (73.21) 37.83 (64.33) 38.27 (44.51) 40.74 (42.13) 35.17 (40.93) 39.81 (43.48)
$c=2/4$ 19.52 (23.17) 29.37 (56.52) 11.28 (35.29) 17.89 (31.79) 11.50 (7.59) 17.89 (14.07)
$c=1/4$ 4.77 (N/A) 13.77 (N/A) 2.22 (N/A) 6.83 (N/A) 0.90 (N/A) 1.87 (N/A)
3 Does Model Invariability Indicate True Knowledge Mastery?
Question 1 offers a potential perspective for evaluating LLMs using MCQA, suggesting that the results on data where the model exhibits invariability may, to some extent, be reliable. Nevertheless, subsequent experiments will challenge this question, and then raise a vital problem within the MCQA-based LLM evaluation through (1) substituting the correct answer options in the MCQs with “None of the above” options and (2) transforming the MCQs into True-or-False format.
3.1 Introducing a “None of the Above” Option
Building upon earlier studies with close-source LLMs [kadavath2022language], we explore the effect of substituting the correct answer option in the MCQs with a “None of the above” option, as illustrated in Figure 2. This modification is applied across the entire selected datasets of MMLU and MedMCQA, as well as on particular subsets (marked as $\blacklozenge$) where LLMs previously responded correctly to all re-ordered answer choices. If the answer to Question 1 is “yes”, in these subsets, LLMs are presumed to possess a genuine comprehension of the knowledge in the questions from the subsets.
As presented in Table 3, as for the experiments on the entire selected datasets, this substitution results in a dramatic decline in LLM performance, with a decrease in both accuracy (Acc) and accuracy-hard (Acc-H) metrics by as much as 70.9%, from 74.44% to a mere 3.54%. Notably, except in a 5-shot scenario involving ChatGPT, the models generally fail to select the “None of the above” option as a substitute instead of the correct answer, with Acc-H falling below 5% in all other scenarios. The performance degradation is even more pronounced within the subsets ($\blacklozenge$), where LLMs, excluding ChatGPT, exhibit even poorer results. Despite prior correct predictions in all re-ordered instances before introducing the “None of the above” option, the results indicate a “No” answer to Question 1, further emphasizing the inadequacies of LLMs in reliably discerning the knowledge encapsulated in MCQA-format questions.
Table 3: Performance deterioration in the LLMs post “None of the Above” substitution. “Acc” denotes the accuracy of LLMs with the initial order of the other options. “Acc-H” represents the proportion of datasets where LLMs answer correctly across all corresponding numbers of re-ordered instances of an original question. $\blacklozenge$ means the subsets of datasets where LLMs have answered correctly across all re-ordered answer options before the alteration of “None of the above” with a baseline accuracy of 100%. Red font delineates the performance decrement compared to that before substitution.